### v2

# IPFormer: Visual 3D Panoptic Scene Completion with Context-Adaptive Instance Proposals

Abstract
[…]
1 Introduction
Figure 1: Comparison of query initialization for Panoptic Scene Completion (PSC). Existing methods, e.g. Symphonies symphonies , randomly initialize instance queries and incorporate context-awareness during training. However, these queries retain their static nature at test time, as they are shared across all inputs. Our method IPFormer initializes them as instance proposals, which preserve context-adaptivity at test time, effectively aggregating directed features for improved PSC performance. Due to the the novelty of vision-based PSC and the absence of established baselines, we apply DBSCAN ester1996_dbscan clustering to Symphonies’ SSC output to retrieve its individual instances.
Panoptic Scene Completion provides a holistic scene understanding that can serve applications like autonomous driving and robotics by reconstructing volumetric geometry from sparse sensor data and assigning meaning to objects in the scene. Such holistic scene understanding requires both accurate geometry and semantic information about a scene. Historically, these tasks have been treated separately and evolved independently due to the distinct nature of geometric reconstruction and semantic interpretation, each requiring specialized algorithms and data representations. Despite the inherent interconnection between geometry and semantics, their separation hindered unified scene understanding. Classical structure-from-motion evolved into simultaneous localization and mapping (SLAM) systems for accurate reconstructions mur2015orb ; engel2014lsd ; koestler2022tandem , while deep learning facilitated monocular depth estimation (MDE) methods godard2017unsupervised ; godard2019digging ; wimbauer2021monorec , which are restricted to visible surfaces. In contrast, Scene Completion refers to the process of inferring the complete 3D structure of a scene, including both visible and occluded regions, from partial observations wimbauer2023behind ; han2024boosting .
Apart from geometry, semantic image segmentation divides an image into semantically meaningful regions and labels them yu2015multi ; chen2017deeplab . Panoptic segmentation kirillov2019panoptic ; kirillov2019panoptic2 ; porzi2019seamless extends this approach by segmenting instances hariharan2014simultaneous for things (countable objects like person or car) while maintaining semantic segmentation for stuff (amorphous regions like road or vegetation). This unified approach is vital for comprehensive scene understanding, as it distinguishes individual object instances while preserving semantic context for background elements, enabling applications like autonomous navigation. The recent trend in scene understanding has been to unify these tasks into Semantic Scene Completion (SSC) hayler2024s4c and Panoptic Scene Completion (PSC) pasco . SSC integrates geometric completion with semantic labeling, predicting both the 3D structure and semantic categories, represented as a 3D grid composed of voxels. PSC builds on this by adding instance-level segmentation for each voxel, thus distinguishing between individual objects. While PSC was introduced using LiDAR modality, methods based on camera images remain largely unexplored. Moreover, recent Transformer-based approaches utilize a fixed set of learned queries to reconstruct objects within the scene volume. Although these queries are typically updated with image context during training, they remain static at test time, limiting their ability to dynamically adapt specifically to the observed scene.
To address these limitations, we introduce IPFormer, a 3D Vision Transformer designed to address Panoptic Scene Completion. IPFormer lifts contextual 2D image features to 3D and samples them based on visibility to initialize instance proposals, each implicitly representing semantic instances within the camera view. Based on this context-adaptive initialization, we establish a robust reconstruction signal that enhances subsequent encoding and decoding stages to reconstruct a complete panoptic scene. Sampling instance proposals solely from visible surfaces is grounded in the principle that all potentially identifiable objects within the observed scene must exhibit a visual cue in the camera image to facilitate completion. This approach of directed feature aggregation at train and test time significantly improves instance identification, semantic classification, and geometric completion.
Our contributions can be summarized as follows:
  • We present IPFormer, the first method that leverages context-adaptive instance proposals at train and test time to address vision-based 3D Panoptic Scene Completion.
  • We introduce a visibility-based sampling strategy, which initializes instance proposals that dynamically adapt to scene characteristics, improving PQ-All by $3.62\text{\,}\%$ and Thing-metrics on average by $18.65\text{\,}\%$, compared to non-adaptive initialization.
[…]
  • Comprehensive ablation studies reveal that employing a dual-head architecture combined with a two-stage training strategy, in which SSC and PSC tasks are trained independently and sequentially, significantly improves performance and effectively balances metrics.
2 Related work
Semantic Scene Completion.   The task of Semantic Scene Completion has seen a large amount of interest after it was first introduced in lmscnet . The initial work focused on solving the task for indoor scenes from camera data zhang2018efficient ; zhang2019cascaded ; liu2018see ; li2019depth ; li2019rgbd ; li2020anisotropic ; chen20203d ; cai2021semantic and used point clouds in outdoor scenarios with LiDAR-based methods roldao20223d ; cheng2021s3cnet ; yan2021sparse ; li2021semisupervised ; rist2021semantic . In the following, we will discuss the most relevant contributions for Semantic Scene Completion in the context of autonomous driving. For a more thorough overview, we refer the reader to the surveys of xu2025survey and zhang2024vision .
Recent progress in perception for autonomous driving has increasingly adopted Vision Transformer architectures dosovitskiyimage , notably DETR carion2020end . DETR introduced learnable object queries, which are trainable embeddings that allow the model to directly predict object sets in parallel, eliminating hand-crafted components like anchor boxes. In SSC, learnable queries can be adapted to predict semantic labels for 3D scene elements (e.g., voxels), streamlining the process and enhancing efficiency. Architectures like OccFormer occformer and VoxFormer li2023voxformer exemplify this trend, with OccFormer using multiscale voxel representations with query features to decode 3D semantic occupancy, and VoxFormer employing depth-based queries fused with image features, refined via deformable attention. Similarly, Symphonies symphonies refines instance queries iteratively using cross-attention between image features and 3D representations, while TPVFormer huang2023tpvformer leverages a tri-perspective view (TPV) to query scene features efficiently. HASSC wang2024not and CGFormer cgformer further refine this general approach by focusing on challenging voxels and contextual queries, respectively. Alternative approaches, such as implicit scenes via self-supervised learning in S4C, hayler2024s4c and CoHFF song2024cohff , which leverages both voxel and TPV representations to fuse features from multiple vehicles, have also emerged, though they do not utilize learnable queries.
It is worth noting that existing literature presents ambiguity regarding context-awareness of queries for SSC. Although both Symphonies symphonies and CGFormer cgformer incorporate contextual cues during training, their approaches diverge at inference. In particular, Symphonies’ instance queries remain static and are shared across all input images, whereas CGFormer’s voxel queries preserve context-awareness by dynamically adapting to each input. To clarify this distinction, we differentiate between instance queries and instance proposals as context-aware and context-adaptive, respectively. Our introduced instance proposals effectively adapt scene characteristics to guide feature aggregation during both train and test time.
Panoptic Scene Completion.   While panoptic LiDAR segmentation, which predicts panoptics for point clouds directly, has been widely studied Che2025S2MC-PHD ; Xiao2024lidarsegopenvoc ; xia2023lidarpanseg , voxel-based PSC approaches have only recently been introduced by PaSCo pasco . This method employs a hybrid convolutional neural network (CNN) and Transformer architecture with static instance queries to address LiDAR-based PSC, enhanced by an uncertainty-aware framework. In parallel, camera-based PSC from multi-frame and multi-view (surround-view) images is addressed by PanoOcc wang2024panoocc , which utilizes static voxel queries to aggregate spatiotemporal information, while SparseOcc liu2023sparseocc introduces a sparse voxel decoder and sparse static instance queries to predict semantic and instance occupancy from up to 96 temporal frames. Concurrently, PanoSSC shi2024panossc introduces a camera-based PSC method using forward-facing, non-temporal imagery. Its dual-head TPV-based architecture separates semantic occupancy and instance segmentation tasks to infer a 3D panoptic voxel grid. Likewise, our method targets PSC using forward-facing, non-temporal imagery. However, it overcomes a key limitation of prior approaches by avoiding reliance on static queries and instead introduces context-adaptive instance proposals that dynamically adapt to the scene at both train and test time. Notably, PanoSSC shi2024panossc is trained on a non-public post-processed version of the SemanticKITTI dataset semantickitti , which, in addition to the absence of released source code, challenges direct comparisons.
3 Methodology
3.1 Overview
Given an input image $\mathbf{X}\in\mathbb{R}^{U\times V\times 3}$ (Fig. 2), the backbone predicts image features $\mathbf{F}\in\mathbb{R}^{H\times W\times F}$, and the depth net predicts a depth map $\mathbf{D}\in\mathbb{R}^{U\times V\times 1}$, following VoxFormer li2023voxformer . Further refining the image features and the depth map allows for generating context-aware 3D representations, denoted as 3D context features $\mathbf{F}_{\text{3D}}$ and initial voxels $\mathbf{V}$. Subsequently, visible and invisible voxels, sampled from $\mathbf{V}$ using the depth map, attend to the 3D image features $\mathbf{F}_{\text{3D}}$ via a series of 3D deformable cross and self-attention mechanisms. This process initializes instance proposals $\mathbf{I}_{P}$ and voxel proposals $\mathbf{V}_{P}$. In this setup, every instance proposal can eventually correlate with a single instance within the observed scene.
The voxel proposals $\mathbf{V}_{P}$ are encoded by 3D local and global encoding, in line with CGFormer cgformer , resulting in voxel features $\mathbf{V}_{F}$. We follow a two-stage training strategy in which the first stage uses these voxel features $\mathbf{V}_{F}$ to address SSC exclusively. This approach effectively guides the latent space towards semantics and geometry. The second training stage targets overall PSC, thereby detecting individual instances. This is achieved by attending the instance proposals $\mathbf{I}_{P}$ over the pre-trained voxel features from the first stage, resulting in instance features $\mathbf{I}_{F}$. These features are used to (i) predict semantics, and (ii) align instances and voxels to aggregate a complete panoptic scene.
Figure 2: Detailed architecture of IPFormer. Our method refines image features and a depth map to produce 3D context features, which are sampled based on visibility to generate context-adaptive instance and voxel proposals. In a two-stage training strategy, voxel proposals first handle Semantic Scene Completion, guiding the latent space toward detailed geometry and semantics. The second stage attends instance proposals over the pretrained voxel features to register individual instances. This dual-head design aligns semantics, instances and voxels, enabling robust Panoptic Scene Completion.
3.2 3D Context Generation
Following CGFormer cgformer , we refine the depth map $\mathbf{D}$ and the feature map $\mathbf{F}\in\mathbb{R}^{H\times W\times F}$ into per-pixel, frustum-shaped depth probabilities $\mathbf{D}_{R}\in\mathbb{R}^{H\times W\times D}$, effectively resulting in $|D|$ depth bins per pixel. This is achieved by applying $\Phi_{D}$ as a mixture of convolution and neighborhood cross-attention layers Hassani2023neigbor , formulated as $\mathbf{D}_{R}=\Phi_{D}(\mathbf{D},\mathbf{F}_{\text{2D}})$. Moreover, the generic image features $\mathbf{F}$ are projected into contextual features $\mathbf{F}_{\text{2D}}\in\mathbb{R}^{H\times W\times C}$ through the context net $\Phi_{C}$, denoted as $\mathbf{F}_{\text{2D}}=\Phi_{C}(\mathbf{F})$.
The next step is to lift context features $\mathbf{F}_{\text{2D}}$ to 3D, following the design of philion2020lift . To elaborate, we lift them to (i) 3D context features $\mathbf{F}_{\text{3D}}$ and (ii) 3D initial voxels $\mathbf{V}$. To achieve this, we distribute $\mathbf{F}_{\text{2D}}$ along rays defined by the camera intrinsics and weighted by the depth probability distribution $\mathbf{D}_{R}$. This is formulated as
$\mathbf{F}_{\text{3D}}(u,v,d,c)=\mathbf{F}_{\text{2D}}(u,v,c)\cdot\mathbf{D}_{R}(u,v,d)\;,$ (1)
where $(u,v)$ are pixel coordinates, $d$ indexes the depth bin, and $c$ is the feature channel dimension. This approach distributes feature vectors across depth bins according to their corresponding probability distributions and thus enables effective lifting of 2D features into a probabilistic 3D volume, while preserving both contextual and spatial information. Note that Eq. (1) does not include learnable parameters.
To additionally create discretized initial voxels $\mathbf{V}\in\mathbb{R}^{X\times Y\times Z\times C}$, we voxelize $\mathbf{F}_{\text{3D}}$ to a regular grid. Let $S=\{(u,v,d)|\mathcal{Q}((u,v,d,c))=(x,y,z)\}$, then
$\mathbf{V}(x,y,z)=\sum_{(u,v,d)\in S}\mathbf{F}_{\text{3D}}(u,v,d)\;,$ (2)
where $\mathcal{Q}$ quantizes continuous coordinates to discrete voxel indices, and $C$ represents the feature dimension. The summation aggregates all features that map to the same voxel, effectively accumulating evidence for stronger feature representations. Building upon the concept of cgformer , we add learnable embeddings to $\mathbf{V}$ that enhance its representational capacity.
3.3 Proposal Initialization
Figure 3: Instance-specific saliency. Through gradient-based attribution, we derive saliency maps that highlight image regions in green, where an individual instance mainly retrieves context from. Our introduced instance proposals effectively adapt to scene characteristics by guiding feature aggregation, substantially improving identification, classification, and completion. In contrast, instance queries sample context in an undirected manner, causing misclassification and geometric ambiguity.
Voxel Visibility.   The initial voxels $\mathbf{V}$ are categorized into visible voxels $\mathbf{V}_{\text{vis}}$ and invisible voxels $\mathbf{V}_{\text{invis}}$, following established SSC works such as li2023voxformer . This is achieved by first applying inverse projection $\Pi^{-1}$ on the depth map $\mathbf{D}$ with camera intrinsics $\mathbf{K}\in\mathbb{R}^{4\times 4}$. This process generates a pseudo point cloud by unprojecting a pixel $(u,v)$ to a corresponding 3D point $(x,y,z)$. Furthermore, let $\mathbf{M}\in\mathbb{R}^{X\times Y\times Z}$ be a binary mask derived from this point cloud that filters the voxel features based on occupancy. Thus, the visible voxel features are defined as $\mathbf{V}_{\text{vis}}=\mathbf{V}\odot\mathbf{M}$, where $\mathbf{M}(x,y,z)=1$ if $(x,y,z)$ corresponds to an unprojected point, and 0 otherwise. The operator $\odot$ denotes the Schur product. Consequently, the invisible voxel features are given by $\mathbf{V}_{\text{invis}}=\mathbf{V}\odot(\mathbf{1}-\mathbf{M})$, with all-ones matrix $\mathbf{1}\in\mathbb{R}^{X\times Y\times Z}$. Note that this sampling does not include learnable parameters.
To further adapt $\mathbf{V}_{\text{vis}}$ with image context, we apply a tailored series of 3D deformable attention mechanisms, inspired by symphonies ; cgformer . Specifically, we apply 3D deformable cross-attention (DCA) to $\mathbf{V}_{\text{vis}}$ (as queries) and the 3D context features $\mathbf{F}_{\text{3D}}$ (as keys and values), resulting in updated visible voxel features $\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}}$. More precisely, an updated voxel query $\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}}$ at 3D location $\mathbf{x}$ computes
$\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}}=\text{DCA}{\left(\mathbf{V}_{\text{vis}},\mathbf{F}_{3D},\mathbf{x}\right)}=\sum_{{n}={1}}^{N}A_{n}W\psi{\left(\mathbf{F}_{3D},\Pi(\mathbf{x})+\Delta{\mathbf{x}}\right)}\;,$ (3)
where $\Pi(\mathbf{x})$ obtains the reference points, $\Delta\mathbf{x}\in\mathbb{R}^{3}$ denotes the estimated displacement from the reference point $\mathbf{x}$, and $\psi(\cdot)$ refers to trilinear interpolation applied to sample from the 3D context features $\mathbf{F}_{\text{3D}}$. The index $n$ loops through the sampled points out of a total of $N$ points, $A_{n}\in[0,1]$ represents the attention weights, and $W$ signifies the transformation weight. We present only the formulation of single-head attention and we utilize multiple layers of deformable cross-attention.
Instance Proposals.   The visible voxels $\tilde{\mathbf{V}}_{\text{vis}}$ are further processed to initialize context-adapted instance proposals $\mathbf{I}_{P}$. Recall from Sec. 1 that instance proposals are initialized at this stage of the architecture, because in principle, all potentially detectable instances within the observed scene must exhibit a visual cue in the camera image to facilitate their completion, thereby defining panoptic scene completion. Following this line of thought, we first apply 3D deformable self-attention (DSA) on $\tilde{\mathbf{V}}_{\text{vis}}$ to foster global context-aware attention over visible voxels. This operation, for a query located at $\mathbf{x}$, is expressed as
$\text{DSA}{(\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}},\tilde{\mathbf{V}}_{\text{vis}},\mathbf{x})}=\sum_{{n}={1}}^{N}A_{n}W\psi{(\tilde{\mathbf{V}}_{\text{vis}},\mathbf{x}+\Delta{\mathbf{x}})}\;.$ (4)
Inspired by CGFormer´s cgformer query generator, our instance generator transforms the DSA output to initialize instance proposals via
$\mathbf{I}_{P}=\text{DSA}(\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}},\tilde{\mathbf{V}}_{\text{vis}},\mathbf{x})+\mathbf{W}_{\mathbf{I}}\in\mathbb{R}^{K\times C}\;,$ (5)
where $\mathbf{W}_{\mathbf{I}}$ represents learnable embeddings to improve the representational capacity, and $K$ denotes the maximum number of detectable instances. This initialization typically directs feature aggregation to highly relevant, instance-dependent image regions, as shown in Fig. 3.
Voxel Proposals.   To initialize voxel proposals $\mathbf{V}_{P}$, we first merge the visible and invisible voxels through element-wise addition along dimension $C$, denoted as $\tilde{\mathbf{V}}_{\text{vis}}\oplus\mathbf{V}_{\text{invis}}$. We then apply 3D deformable self-attention to distribute updated context information over the entire scene volume, formulated as $\mathbf{V}_{P}=\text{DSA}(\tilde{\mathbf{V}}_{\text{vis}}\oplus\mathbf{V}_{\text{invis}})$, which is in line with li2023voxformer ; symphonies ; cgformer :
3.4 Encoding
The purpose of this part of our architecture is to (i) transform voxel proposals $\mathbf{V}_{P}$ into voxel features $\mathbf{V}_{F}$ that encode semantic information and (ii) project this semantic information onto the instance proposals $\mathbf{I}_{P}$, thus prompting semantic instance-voxel relationships.
To encode semantic information, we propagate the voxel proposals $\mathbf{V}_{P}$ through a 3D Local and Global Encoder (LGE) based on CGFormer cgformer . This encoder enhances the semantic and geometric discriminability of the scene volume by integrating TPV huang2023tpvformer and voxel representations, effectively capturing global and local contextual features, respectively. For an in-depth discussion, we direct readers to cgformer . The 3D LGE $\Phi_{\text{LGE}}$ transforms voxel proposals $\mathbf{V}_{P}$ into semantic voxel features $\mathbf{V}_{F}\in\mathbb{R}^{X\times Y\times Z\times C}$, formulated as $\mathbf{V}_{F}=\Phi_{\text{LGE}}(\mathbf{V}_{P})$. Building on these semantically encoded voxel features, we apply cross-attention (CA) to instance proposals $\mathbf{I}_{P}\in\mathbb{R}^{K\times C}$ (as queries) and semantic voxel features $\mathbf{V}_{F}\in\mathbb{R}^{X\times Y\times Z\times C}$ (as keys and values), producing updated instance features $\tilde{\mathbf{I}}_{P}^{\text{CA}}$. Specifically, the updated instance features are computed as:
$\tilde{\mathbf{I}}_{P}^{\text{CA}}=\text{CA}(\mathbf{I}_{P},\mathbf{V}_{F})=\sum_{n=1}^{N}A_{n}W\mathbf{V}_{F}(\mathbf{p}_{n})\;,$ (6)
where $N=X\cdot Y\cdot Z$ denotes the total number of voxels, $\mathbf{p}_{n}$ represents the $n$-th voxel position, $A_{n}\in[0,1]$ is the attention weight, and $W$ is the learnable projection weight.
Subsequently, we apply self-attention (SA) to the updated instance features $\tilde{\mathbf{I}}_{P}^{\text{CA}}$ to capture global dependencies among the $K$ instances, yielding further refined instances $\tilde{\mathbf{I}}_{P}^{\text{SA}}$. The self-attention mechanism is defined as:
$\tilde{\mathbf{I}}_{P}^{\text{SA}}=\text{SA}(\tilde{\mathbf{I}}_{P}^{\text{CA}})=\sum_{n=1}^{K}A_{n}W\tilde{\mathbf{I}}_{P}^{\text{CA}}(n)\;,$ (7)
where $\tilde{\mathbf{I}}_{P}^{\text{CA}}(n)$ denotes the feature vector of the $n$-th instance from $\tilde{\mathbf{I}}_{P}^{\text{CA}}$. To obtain the final encoded semantic instance features, we further process the self-attention output $\tilde{\mathbf{I}}_{P}^{\text{SA}}$ through an MLP Encoder $\Phi_{E}$, yielding $\mathbf{I}_{F}\in\mathbb{R}^{K\times C}$. This step refines the instance representations for downstream tasks. The encoding is computed as $\mathbf{I}_{F}=\Phi_{E}(\tilde{\mathbf{I}}_{P}^{\text{SA}})\in\mathbb{R}^{K\times C}$.
3.5 Decoding
Since we employ a two-stage training scheme (see Sec. 3.6), the goal of decoding is twofold. In the first stage, a lightweight MLP head $\phi_{\text{SSC}}$ predicts the semantic scene via $\phi_{\text{SSC}}(\mathbf{V}_{F})$.
In the second training stage, a lightweight panoptic head decodes instance features $\mathbf{I}_{F}$ to semantic class logits $\mathbf{I}_{L}\in\mathbb{R}^{K\times L}$, where $K$ is the number of instances and $L$ is the number of semantic classes. These logits are further processed with voxel features $\mathbf{V}_{F}$ to perform instance-voxel alignment that eventually enables panoptic scene aggregation, inspired by PaSCo pasco . To elaborate, we align the instance features $\mathbf{I}_{F}$ with the voxel features $\mathbf{V}_{F}\in\mathbb{R}^{X\times Y\times Z\times C}$ to compute affinity scores $\mathbf{H}\in\mathbb{R}^{X\times Y\times Z\times K}$ by applying the dot product, resulting in $\mathbf{H}=\mathbf{V}_{F}\cdot\mathbf{I}_{F}^{\top}$. These scores are converted to predicted probabilities via $\mathbf{H}_{P}=\text{sigmoid}(\mathbf{H})$, where $\mathbf{H}_{P}$ denotes the model’s confidence in each voxel belonging to each instance. For panoptic aggregation pasco , voxel-wise logits $\mathbf{V}_{L}\in\mathbb{R}^{X\times Y\times Z\times L}$ are computed as $\mathbf{V}_{L}=\mathbf{H}_{P}\cdot\mathbf{I}_{L}$. It should be emphasized that voxel-wise logits do not represent the probability of belonging to a certain semantic class, but the confidence of belonging to one of the instances. Furthermore, semantic class IDs are assigned via argmax over $\mathbf{V}_{L}$ along $L$, and instance IDs via argmax over $\mathbf{H}_{P}^{\top}$ along $K$, yielding a complete panoptic scene. Note that the alignment process and the panoptic aggregation do not include learnable parameters.
3.6 Training Objective
We follow a two-stage training strategy in which we formulate the learning objectives for SSC and PSC tasks sequentially. Technically, the first stage optimizes
$\mathcal{L}_{\text{SSC}}=\lambda_{\text{ce}}\mathcal{L}_{\text{ce}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{sem}}\mathcal{L}_{\text{scal}}^{\text{sem}}+\lambda_{\text{geo}}\mathcal{L}_{\text{scal}}^{\text{geo}}\;,$ (8)
where the weights for the cross-entropy loss, semantic scene-class affinity loss (SCAL), and geometric SCAL monoscene are set to $\lambda_{\text{ce}}=\lambda_{\text{sem}}=\lambda_{\text{geo}}=1$, respectively, and the weight for the depth loss cgformer is set to $\lambda_{\text{depth}}=10^{-4}$.
The second stage calculates losses mask-wise by comparing each ground-truth mask with the best matching predicted mask. Consistent with previous research pasco ; shi2024panossc , we find the best match by applying the Hungarian method kuhn1955hungarian to perform bipartite matching carion2020end , using an IoU threshold of $50\text{\%}$. This threshold follows the standard convention established in prior work on panoptic segmentation kirillov2019panoptic ; pasco , and the Hungarian algorithm ensures a globally optimal, permutation-invariant one-to-one assignment. After matching, the second stage optimizes
$\mathcal{L}_{\text{PSC}}=\lambda_{\text{mask}}\mathcal{L}_{\text{ce}}^{\text{mask}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}}+\lambda_{\text{focal}}\mathcal{L}_{\text{focal}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}\;,$ (9)
where $\mathcal{L}_{\text{ce}}^{\text{mask}}$ represents mask-wise cross-entropy loss, $\mathcal{L}_{dice}$ is the dice loss of Milletari_2016_diceloss_vnet and $\mathcal{L}_{focal}$ is the focal loss based on Lin_2017_focalloss . In practice, we set the weights of these losses to $\lambda_{\text{mask}}=\lambda_{\text{dice}}=1$ and $\lambda_{\text{focal}}=40$, respectively, following PaSco pasco . Moreover, $\mathcal{L}_{\text{depth}}$ is defined identically to that in $\mathcal{L}_{\text{SSC}}$.
4 Experiments
4.1 Quantitative Results
Dataset and Baselines.   We conduct our experiments by (i) in-domain training and testing on the SemanticKITTI SSC dataset semantickitti , and (ii) out-of-domain zero-shot generalization on the distinct SSCBench-KITTI360 li2024sscbench . The instance ground-truths for both datasets are provided by PaSCo pasco . Given the novelty of vision-based PSC, no established baselines currently exist for direct comparison. Furthermore, recall from Sec. 2 that PanoSSC shi2024panossc represents an existing vision-based PSC method using forward-facing imagery. This would make it a suitable baseline for comparison. However, this approach is trained on a non-public post-processed version of the SemanticKITTI dataset semantickitti , which, in addition to the absence of released source code, challenges direct comparison with our method (see Sec. A.6 of the technical appendix for further details). Consequently, we adapt the latest state-of-the-art vision-based SSC methods monoscene ; symphonies ; occformer ; cgformer and generate instance predictions by clustering Thing-classes from their outputs, in line with PaSCo pasco , which introduced the task of 3D PSC. In particular, to ensure fairness and methodological rigor, we apply DBSCAN ester1996_dbscan , the same Euclidean clustering algorithm used to construct the PSC ground truth in PaSCo. Furthermore, we utilize pre-trained checkpoints for the baselines, acquired from the official publicly available implementations. We provide further details in Sec. A.1 of the technical appendix.
In-Domain Performance.   In summary, IPFormer exceeds all baselines in overall panoptic metrics PQ and PQ, and achieves best or second-best results on individual metrics, as shown in Tab. 1. These results showcase the significant advancements IPFormer brings to Panoptic Scene Completion, highlighting its robustness and efficiency. Moreover, IPFormer attains superior performance on RQ-All, while securing the second-best position in SQ-All and PQ-Thing. Furthermore, although our approach places second in SQ-Stuff, it surpasses existing methods in RQ-Stuff and notably PQ-Stuff, indicating superior capability in recognizing Stuff-classes accurately. Notably, despite our method achieving state-of-the-art performance in PSC, it exhibits moderate SSC performance on in-domain data. These results are evaluated on the PSC output of our two-staged, fully-trained model. For SSC metrics, the instance IDs of all voxels are disregarded, effectively reducing PSC to SSC. Additionally, our method directly predicts a full panoptic scene, resulting in a significantly superior runtime of $0.33\text{\,}\mathrm{s}$, thus providing a runtime reduction of over $14\times$. In the technical appendix, we further report class-wise quantitative results and a detailed runtime analysis, including memory utilization during training.
Out-of-Domain Zero-Shot Generalization Performance.   Trained on SemanticKITTI semantickitti , we cross-validate our method on the distinct SSCBench-KITTI360 dataset li2024sscbench . Results in Tab. 2 show that IPFormer demonstrates superior zero-shot generalization capability. Especially in comparison to its closest baseline, CGFormer+DBSCAN, which it outperforms across PSC and SSC metrics. These results highlight the generalization capability of IPFormer by leveraging out-of-domain image context to robustly initialize context-adaptive instance proposals, resulting in superior SSC and PSC performance.
Table 1: In-domain performance on SemanticKITTI val. set semantickitti . Best and second-best results are bold and underlined, respectively. Due to the absence of established baselines for vision-based PSC (see Sec. 4.1), we infer state-of-the-art SSC methods and apply DBSCAN ester1996_dbscan to retrieve instances.
    PSC Metrics     SSC Metrics
    All Thing Stuff
[…]
MonoScene monoscene + DBSCAN     $10.12$ $3.43$ $15.15$ $5.33$ $0.51$ $7.36$ $0.87$ $5.56$ $20.81$ $8.57$     $36.80$ $11.31$     $\underline{4.51}$
Symphonies symphonies + DBSCAN     $11.69$ $3.75$ $26.09$ $5.95$ $1.07$ $27.65$ $1.76$ $5.70$ $24.95$ $8.99$     $\underline{41.92}$ $15.02$     $4.54$
OccFormer occformer + DBSCAN     $11.25$ $4.32$ $24.19$ $6.69$ $0.68$ $21.47$ $1.15$ $6.96$ $26.16$ $10.73$     $36.43$ $13.51$     $4.70$
CGFormer cgformer + DBSCAN     $\underline{14.39}$ $\underline{6.16}$ $\mathbf{48.14}$ $\underline{9.48}$ $\mathbf{2.20}$ $\mathbf{44.46}$ $\mathbf{3.47}$ $\underline{9.03}$ $\mathbf{50.82}$ $\underline{13.86}$     $\mathbf{45.98}$ $\mathbf{16.89}$     $4.70$
IPFormer (ours)     $\mathbf{14.45}$ $\mathbf{6.30}$ $\underline{41.95}$ $\mathbf{9.75}$ $\underline{2.09}$ $\underline{42.67}$ $\underline{3.33}$ $\mathbf{9.35}$ $\underline{41.43}$ $\mathbf{14.43}$     $40.90$ $\underline{15.33}$     $\mathbf{0.33}$
Table 2: Out-of-domain zero-shot generalization performance of IPFormer and the closest baseline CGFormer+DBSCAN, by training on SemanticKITTI semantickitti and cross-validating on SSCBench-KITTI360 test set li2024sscbench . IPFormer demonstrates superior absolute and relative generalization performance across PSC and SSC metrics.
    PSC Metrics     SSC Metrics
    All Thing Stuff
[…]
SemanticKITTI
CGFormer cgformer + DBSCAN     $14.39$ $6.16$ $\mathbf{48.14}$ $9.48$ $\mathbf{2.20}$ $\mathbf{44.46}$ $\mathbf{3.47}$ $9.03$ $\mathbf{50.82}$ $13.86$     $\mathbf{45.98}$ $\mathbf{16.89}$
IPFormer (ours)     $\mathbf{14.45}$ $\mathbf{6.30}$ $41.95$ $\mathbf{9.75}$ $2.09$ $42.67$ $3.33$ $\mathbf{9.35}$ $41.43$ $\mathbf{14.43}$     $40.90$ $15.33$
KITTI-360
CGFormer cgformer + DBSCAN     $8.44$ $1.08$ $17.82$ $1.87$ $\mathbf{0.53}$ $20.06$ $\mathbf{0.96}$ $1.48$ $16.19$ $2.54$     $28.11$ $9.44$
IPFormer (ours)     $\mathbf{9.41}$ $\mathbf{1.23}$ $\mathbf{24.68}$ $\mathbf{2.16}$ $0.52$ $\mathbf{22.76}$ $0.95$ $\mathbf{1.68}$ $\mathbf{25.89}$ $\mathbf{2.93}$     $\mathbf{28.74}$ $\mathbf{9.53}$
Relative Gap $\downarrow$
CGFormer cgformer + DBSCAN     $41.37\%$ $82.47\%$ $62.98\%$ $80.28\%$ $75.91\%$ $54.89\%$ $72.34\%$ $83.61\%$ $68.15\%$ $81.67\%$     $38.88\%$ $44.09\%$
IPFormer (ours)     34.88% 80.48% 41.19% 77.85% 75.12% 46.64% 71.53% 82.03% 37.52% 79.69%     29.73% 37.81%
Table 3: Ablation on the initialization of instance queries vs. instance proposals.
instance instance
Method queries proposals
All
[…]
Thing
[…]
Stuff
[…]
Table 4: Ablation on the visibility-based sampling strategy for proposal initialization.
visible + visible
Method invisible only
All
[…]
Thing
[…]
Stuff
[…]
Table 5: Ablation on deep supervision for instance encoding during second-stage training.
w/ deep w/o deep
Method supervision supervision
All
[…]
Thing
[…]
Stuff
[…]
Table 6: Ablation on the dual-head design and the training strategy. Methods (a)-(d) evaluate combinations of single/dual-head and one/two-stage trainings, where $\ast$ denotes frozen weights of the first stage during second-stage training. Methods (e)-(i) examine the incorporation of the first-stage loss $\mathcal{L}_{\text{SSC}}$ by different factors of $\lambda_{\text{SSC}}$ into the second stage. (j) represents the final IPFormer config.
Dual Two All Thing Stuff
[…]
(a) - $\underline{14.64}$ $6.23$ $33.37$ $9.54$ $1.75$ $15.40$ $2.66$ $\mathbf{9.49}$ $\mathbf{46.44}$ $\mathbf{14.55}$
(b) - $14.21$ $5.57$ $36.09$ 8.58 $1.06$ $22.29$ $1.73$ $8.84$ $\underline{46.12}$ $13.57$
(c) - $10.66$ $0.42$ $14.31$ $0.76$ $0.13$ $7.07$ $0.22$ $0.63$ $19.57$ $1.15$
(d)   $\ast$ - $13.75$ $5.06$ $38.15$ $7.98$ $0.27$ $27.31$ $0.49$ $8.55$ $46.00$ $13.43$
(e) $1.00$ $14.35$ $6.06$ $\underline{38.74}$ $9.30$ $1.66$ $35.52$ $2.63$ $9.27$ $41.08$ $14.16$
(f) $0.50$ $14.42$ $\underline{6.27}$ $38.72$ $\underline{9.69}$ $\underline{1.88}$ $\underline{42.01}$ $\underline{3.13}$ $\underline{9.46}$ $36.33$ $\mathbf{14.45}$
(g) $0.20$ $14.46$ $6.05$ $33.06$ $9.37$ $1.60$ $28.79$ $2.54$ $9.29$ $36.18$ $14.34$
(h) $0.10$ $14.41$ $6.08$ $33.50$ $9.40$ $1.72$ $29.93$ $2.73$ $9.26$ $36.27$ $14.26$
(i) $0.01$ $\mathbf{14.67}$ $5.16$ $24.33$ $7.97$ $0.90$ $7.83$ $1.43$ $8.27$ $36.32$ $12.73$
(j) - $14.45$ $\mathbf{6.30}$ $\mathbf{41.95}$ $\mathbf{9.75}$ $\mathbf{2.09}$ $\mathbf{42.67}$ $\mathbf{3.33}$ $9.35$ $41.43$ $14.43$
4.2 Ablation Study
Context-Adaptive Initialization.  We compare our introduced context-adaptive instance proposals with context-aware, but non-adaptive instance queries (Tab. 5). While initializing proposals adaptively slightly decreases SQ-Stuff, it remarkably increases all Thing-related metrics by $18.65\text{\,}\%$ on average, with RQ-Things showing the best performance gain of $21.98\text{\,}\%$. This insight demonstrates that adaptively initializing instance proposals from image context not only enables us to drastically recognize more objects, but also enables Thing segmentation with substantially higher quality.
Visibility-Based Sampling Strategy.  We examine our proposed visible-only approach for proposal initialization to a method that employs both visible and invisible voxels (Tab. 5). Formally, these proposals are initialized similar to Eq. (5), specifically as $\text{DSA}(\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}}\oplus\mathbf{V}_{\text{invis}}^{\mathbf{x}},\tilde{\mathbf{V}}_{\text{vis}}\oplus\mathbf{V}_{\text{invis}},\mathbf{x})+\mathbf{W}_{\mathbf{I}}$. Under our visible-only approach, almost all metrics increase significantly, most notably RQ-Thing and SQ-Thing by $48.00\text{\,}\%$ and $93.95\text{\,}\%$, respectively, thus evidently enhancing recognition and segmentation of objects from the Thing-category. These findings confirm that visible voxels carry a robust reconstruction signal, which can be leveraged to initialize instance proposals and eventually improve PSC performance.
Deep Supervision.  Inspired by PanoSSC shi2024panossc , we investigate deep supervision in the form of auxiliary losses, essentially supervising the attention maps of each layer during the encoding of the instance proposals (Tab. 5). Our findings show that guiding intermediate layers in such way degrades all metrics. Most notably, when deep supervision is not applied, SQ-Thing registers a substantial improvement of $102.61\text{\,}\%$.
Dual-head and Two-Stage Training.  The baseline (a) in Tab. 6 uses a single head and single-stage training, performing well on Stuff but poorly in SQ-Thing. Introducing separate heads for SSC and PSC in (b) slightly reduces Stuff performance but improves SQ-Thing to nearly match the baseline. Replacing the dual-head with a purely two-stage approach in (c) severely degrades SQ and yields the worst RQ scores. In contrast, adopting both the dual-head architecture and two-stage training in (j) achieves the best overall results. Alternatively, freezing the first stage during the second-stage training (d) proves detrimental, notably reducing RQ-Thing. Furthermore, methods (e)-(i) evaluate the integration of the SSC objective into the second-stage training, by adding $\mathcal{L}_{\text{SSC}}$ to $\mathcal{L}_{\text{PSC}}$ via different factors of $\lambda_{\text{SSC}}$. Overall, these configurations fall short of the superior performance achieved by our final method (j), which applies both a dual-head architecture and a two-stage training strategy that separates SSC and PSC objectives. Note that these findings are underscored by additional ablation experiments presented in Sec. A.7 of the technical appendix.
Figure 4: Qualitative results on the SemanticKITTI val. set semantickitti . Each top row illustrates purely semantic information, following the SSC color map. Each bottom row displays individual instances, with randomly assigned colors to facilitate differentiation. Note that we specifically show instances of the Thing-category for clarity.
4.3 Qualitative Results
Presented in Tab. 4, IPFormer surpasses existing approaches by excelling at identifying individual instances, inferring their semantics, and reconstructing geometry with exceptional fidelity. Even for extremely low-frequency categories such as the person category ($0.07\text{\,}\%$) under adverse lighting conditions, and in the presence of trace-artifacts from dynamic objects in the ground-truth data, our method proves visually superior. These advancements stem from IPFormer’s instance proposals, which dynamically adapt to scene characteristics, thus preserving high precision in instance identification, semantic segmentation, and geometric completion. Conversely, other models tend to encounter challenges in identifying semantic instances effectively while simultaneously retaining geometric integrity. Moreover, our instance-specific saliency analysis in Fig. 3 underscores these findings.
5 Conclusion
[…]
Appendix A Technical Appendix
A.1 Experimental setup
[…]
To enable panoptic evaluation, we adopt the PaSCo dataset [3], which extends by generating pseudo panoptic labels. PaSCo employs DBSCAN [13] to cluster voxels of “thing” classes into distinct instance IDs, using a distance threshold of $\epsilon=1$ and a minimum group size of $\text{MinPts}=8$. The authors of PaSCo ensure that the pseudo labels are valid by comparing them against the available LiDAR single-scan point-wise panoptic ground truth from the validation set. Both the pseudo labels (generated by DBSCAN) and the ground truth are voxelized, and their quality is assessed in regions where both are defined. For more details, including quantitative and qualitative evaluation, see PaSco [3, supplementary material Sec. 8.2]. Since instance labels cannot be derived for SemanticKITTI’s hidden test set, we perform evaluations on the validation set.
Metrics.   To evaluate our Panoptic Scene Completion approach, we employ the Panoptic Quality (PQ) metric [24], which combines Segmentation Quality (SQ) and Recognition Quality (RQ). PQ is defined as:
$\text{PQ}=\text{SQ}\times\text{RQ}=\frac{\sum_{(p,g)\in TP}\text{IoU}(p,g)}{|TP|+\frac{1}{2}|FP|+\frac{1}{2}|FN|}\;,$ (10)
where $\text{SQ}=\frac{\sum_{(p,g)\in TP}\text{IoU}(p,g)}{|TP|}$ and $\text{RQ}=\frac{|TP|}{|TP|+\frac{1}{2}|FP|+\frac{1}{2}|FN|}$. $TP$, FPFP, and $FN$ represent true positives, false positives, and false negatives, respectively, and IoU is the intersection-over-union. SQ measures segmentation fidelity via the average IoU of matched segments, while RQ assesses recognition accuracy as the F1-score [10]. We compute PQ, SQ, and RQ across all classes, as well as separately for “stuff” (amorphous regions) and “things” (countable objects), to analyze category-specific performance.
The standard PQ metric requires a predicted segment to match a ground-truth segment with $\text{IoU}>0.5$. However, this strict threshold can be overly conservative for Stuff classes, which typically lack well-defined boundaries. The metric PQ [42] relaxes the matching criterion specifically for Stuff classes. Formally, $\text{PQ}^{\dagger}$ retains the same formulation as PQ:
$\text{PQ}^{\dagger}=\frac{\sum_{(p,g)\in\text{TP}^{\dagger}}\text{IoU}(p,g)}{|\text{TP}^{\dagger}|+\frac{1}{2}|\text{FP}^{\dagger}|+\frac{1}{2}|\text{FN}^{\dagger}|},$ (11)
but relaxes the matching condition used to define true positives ($\text{TP}^{\dagger}$), and thus false positives ($\text{FP}^{\dagger}$) and false negatives ($\text{FN}^{\dagger}$). Specifically, for Thing classes, predicted and ground-truth segment pairs $(p,g)$ are matched if $\text{IoU}(p,g)>0.5$, identical to the original PQ definition. For Stuff classes, matches are accepted if $\text{IoU}(p,g)>0$, thereby allowing any overlapping prediction to contribute to the metric. This relaxation acknowledges the inherent ambiguity in delineating stuff regions and reduces penalties for minor misalignments. As with PQ, we compute PQ jointly across all classes and separately for stuff and thing categories to enable detailed performance analysis.
A.2 Implementation Details
Training and Architecture.
In accordance with [4, 20, 32, 21], we train for $25$ epochs in the first stage and $30$ epochs in the second stage, using AdamW [36] optimizer with standard hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a batch size of $1$. We utilize a single NVIDIA A100 80GB GPU, adopt a maximum learning rate of $1\times 10^{-4}$, and implement a cosine adaptive learning rate schedule decay, with a cosine warmup applied over the initial $2$ epochs. Our implementation is based on PyTorch [39] with an fp32 backend. Moreover, we operate on a $50\text{\,}\%$ voxel grid resolution of $X=128$, $Y=128$, $Z=16$ and finally upsample to the ground-truth grid resolution of $256\times 256\times 32$ via trilinear interpolation. The feature dimension is set to $C=128$. Training IPFormer takes approximately $3.5$ days for each of the two stages. The second stage training is initialized with the final model state of the first stage, and we eventually present results for the best checkpoint based on $\text{PQ}^{{\dagger}}$. Aligning with [32, 21, 59], we adopt a pretrained MobileStereoNet [46] to estimate depth maps, and employ EfficientNetB7 [49] as our image backbone, consistent with [63, 59]. Moreover, the context net consists of a lightweight CNN, while the panoptic head represents a single linear layer for projection to class logits. The deformable cross and self attention blocks during proposal initialization consist of three layers and two layers, respectively, while 8 points are sampled for each reference point. Finally, the cross and self-attention blocks during decoding each consist of three layers.
Clustering.
To cluster the predictions of SSC baselines and retrieve individual instances, we apply DBSCAN [13] with parameters $\epsilon=1$ and MinPts=8, in line with the work of PaSCo [3], which provides ground-truth instances for the SemanticKITTI dataset [1]. The clustering is performed on an AMD EPYC 7713 CPU (allocating 16 cores) with 64GB memory.
A.3 Class-Wise Quantitative Results
In addition to the overall performance on Panoptic Scene Completion in Tab. 1, we report class-wise results in Tab. 7. IPFormer consistently ranks first or second, thereby demonstrating state-of-the-art performance in vision-based PSC, aligning with our primary findings.
As also shown in Tab. 1, all methods showcase suboptimal performance on Thing classes, which arises from the significant class imbalance in the SemanticKITTI dataset, where Thing classes make up only $4.53\text{\,}\%$ of all voxels. To address this, we employ the Sigmoid Focal Loss (Eq. 12), which down-weights easier examples and focuses on harder-to-classify ones, particularly rare Thing classes. Sec. A.7 presents and discusses additional ablation results on the Focal Loss to demonstrate its effectiveness. Our proposed method balances performance between Thing and Stuff classes, achieving state-of-the-art results by prioritizing equitable performance across both categories rather than solely optimizing for Thing classes.
Table 7: Class-wise quantitative results on SemanticKITTI val. set [1] (best, second-best) with corresponding class frequencies. The asterisk ($\ast$) indicates SSC methods, for which the outputs are clustered to identify their instances, as described in Sec. 4.1.
Method
$\blacksquare$ Car (3.92%)
$\blacksquare$ Bicycle (0.03%)
$\blacksquare$ Motorcycle (0.03%)
$\blacksquare$ Truck (0.16%)
$\blacksquare$ Other-Vehicle  (0.20%)
$\blacksquare$ Person (0.07%)
$\blacksquare$ Bicyclist (0.07%)
$\blacksquare$ Motorcyclist. (0.05%)
$\blacksquare$ Road (15.30%)
$\blacksquare$ Parking (1.12%)
$\blacksquare$ Sidewalk (11.13%)
$\blacksquare$ Other-Ground (0.56%)
$\blacksquare$ Building (14.10%)
$\blacksquare$ Fence (3.90%)
$\blacksquare$ Vegetation (39.30%)
$\blacksquare$ Trunk (0.51%)
$\blacksquare$ Terrain (9.17%)
$\blacksquare$ Pole (0.29%)
$\blacksquare$ Traffic-Sign (0.08%)
Mean
$\mathbf{PQ}$ MonoScene [4]* $4.10$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $53.92$ $1.27$ $1.01$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $4.95$ $0.00$ $0.00$ $3.43$
Symphonies [21]* $7.80$ $0.00$ $0.00$ $0.14$ $0.37$ $0.22$ $0.00$ $0.00$ $54.79$ $0.46$ $1.92$ $0.00$ $0.00$ $0.00$ $0.63$ $0.00$ $4.88$ $0.00$ $0.00$ $3.75$
OccFormer [63]* $3.51$ $0.00$ $0.00$ $\underline{1.60}$ $0.33$ $0.00$ $0.00$ $0.00$ $59.52$ $\underline{3.49}$ $6.79$ $0.00$ $0.00$ $0.00$ $0.51$ $0.00$ $6.30$ $0.00$ $0.00$ $4.32$
CGFormer [59]* $\mathbf{14.14}$ $\mathbf{0.58}$ $\mathbf{1.14}$ $0.99$ $\underline{0.51}$ $\underline{0.25}$ $0.00$ $0.00$ $\mathbf{66.78}$ $3.43$ $\underline{11.08}$ $0.00$ $\mathbf{0.48}$ $\mathbf{0.09}$ $\underline{1.14}$ $\mathbf{0.10}$ $\mathbf{15.35}$ $\mathbf{0.63}$ $\mathbf{0.28}$ $\underline{6.16}$
IPFormer (ours) $\underline{12.83}$ $\underline{0.45}$ $\underline{0.56}$ $\mathbf{1.69}$ $\mathbf{0.92}$ $\mathbf{0.27}$ $0.00$ $0.00$ $\underline{66.28}$ $\mathbf{6.16}$ $\mathbf{14.52}$ $0.00$ $\underline{0.13}$ $0.00$ $\mathbf{2.28}$ $0.00$ $\underline{13.25}$ $\underline{0.12}$ $\underline{0.15}$ $\mathbf{6.30}$
$\mathbf{SQ}$ Monoscene [4]* $58.89$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $66.55$ $55.27$ $52.11$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $55.00$ $0.00$ $0.00$ $15.15$
Symphonies [21]* $61.45$ $0.00$ $0.00$ $51.35$ $52.88$ $55.47$ $0.00$ $0.00$ $65.11$ $50.71$ $53.09$ $0.00$ $0.00$ $0.00$ $51.64$ $0.00$ $53.94$ $0.00$ $0.00$ $26.09$
OccFormer [63]* $58.62$ $0.00$ $0.00$ $\mathbf{61.53}$ $51.59$ $0.00$ $0.00$ $0.00$ $68.26$ $\underline{56.29}$ $54.69$ $0.00$ $0.00$ $0.00$ $52.22$ $0.00$ $56.33$ $0.00$ $0.00$ $24.19$
CGFormer [59]* $\underline{65.59}$ $\mathbf{52.01}$ $\underline{54.16}$ $\underline{57.16}$ $\mathbf{59.22}$ $\mathbf{67.53}$ $0.00$ $0.00$ $\underline{70.37}$ $54.51$ $\underline{55.15}$ $0.00$ $\underline{50.91}$ $\mathbf{52.86}$ $\underline{52.41}$ $\mathbf{51.43}$ $\underline{58.94}$ $\mathbf{58.72}$ $\mathbf{53.72}$ $\mathbf{48.14}$
IPFormer (ours) $\mathbf{65.84}$ $\underline{51.25}$ $\mathbf{58.60}$ $52.34$ $\underline{57.34}$ $\underline{56.02}$ $0.00$ $0.00$ $\mathbf{70.40}$ $\mathbf{56.44}$ $\mathbf{55.36}$ $0.00$ $\mathbf{52.93}$ $0.00$ $\mathbf{52.68}$ $0.00$ $\mathbf{59.37}$ $\underline{57.59}$ $\underline{50.91}$ $\underline{41.95}$
$\mathbf{RQ}$ MonoScene[4]* $6.97$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $81.02$ $2.30$ $1.94$ $0.00$ $0.00$ $0.00$ $0.00$ $0.00$ $9.01$ $0.00$ $0.00$ $5.33$
Symphonies[21]* $12.69$ $0.00$ $0.00$ $0.27$ $0.70$ $\underline{0.40}$ $0.00$ $0.00$ $84.15$ $0.91$ $3.61$ $0.00$ $0.00$ $0.00$ $1.22$ $0.00$ $9.04$ $0.00$ $0.00$ $5.95$
OccFormer[63]* $5.98$ $0.00$ $0.00$ $\underline{2.60}$ $0.64$ $0.00$ $0.00$ $0.00$ $87.20$ $6.20$ $12.41$ $0.00$ $0.00$ $0.00$ $0.98$ $0.00$ $11.19$ $0.00$ $0.00$ $6.69$
CGFormer[59]* $\mathbf{21.56}$ $\mathbf{1.11}$ $\mathbf{2.11}$ $1.73$ $\underline{0.87}$ $0.37$ $0.00$ $0.00$ $\mathbf{94.91}$ $\underline{6.30}$ $\underline{20.09}$ $0.00$ $\mathbf{0.94}$ $\mathbf{0.17}$ $\underline{2.18}$ $\mathbf{0.20}$ $\mathbf{26.05}$ $\mathbf{1.06}$ $\mathbf{0.51}$ $\underline{9.48}$
IPFormer (ours) $\underline{19.49}$ $\underline{0.88}$ $\underline{0.95}$ $\mathbf{3.23}$ $\mathbf{1.60}$ $\mathbf{0.49}$ $0.00$ $0.00$ $\underline{94.16}$ $\mathbf{10.91}$ $\mathbf{26.23}$ $0.00$ $\underline{0.24}$ $0.00$ $\mathbf{4.32}$ $0.00$ $\underline{22.31}$ $\underline{0.22}$ $\underline{0.30}$ $\mathbf{9.75}$
A.4 Additional Qualitative Results.
We provide further qualitative results in Fig. 5, aligning with the primary results in that IPFormer reconstructs and identifies diverse objects of various sizes.
A.5 Compute Resources and Runtime
In Tab. 8, we provide memory utilization during training and a detailed runtime analysis, in combination with the resulting performance for PQ. Our method has the highest memory utilization with $52.8\text{\,}\mathrm{G}\mathrm{B}$, while OccFormer has the lowest with $17.3\text{\,}\mathrm{G}\mathrm{B}$. As elaborated in Sec. 4.1, we retrieve PSC predictions for the baselines by clustering their SSC predictions. Consequently, the total runtime for all baselines consists of inference time in addition to clustering time, with the latter being a constant of $4.43\text{\,}\mathrm{s}$ seconds for all baselines. Since our method directly predicts a panoptic scene, the total runtime is equal to the inference time. Thus, IPFormer exhibits a significantly superior runtime of $0.33\text{\,}\mathrm{s}$ compared to the second-best method in terms of PQ, CGFormer, which has a total runtime of $4.70\text{\,}\mathrm{s}$. IPFormer therefore provides a runtime reduction of over $14\times$.
Figure 5: Additional qualitative results on the SemanticKITTI val. set [1]. Each top row illustrates purely semantic information, following the SSC color map. Each bottom row displays individual instances, with randomly assigned colors to facilitate differentiation. Note that we specifically show instances of the Thing category for clarity.
Table 8: Comparison of compute resources during training, and detailed runtime analysis. Additionally, we show the PQ metric on SemanticKITTI val. set [1] (best, second-best). Operations are performed on a single NVIDIA A100 GPU with $80\text{\,}\mathrm{G}\mathrm{B}$ memory and an AMD EPYC 7713 CPU (allocating 16 cores) with $64\text{\,}\mathrm{G}\mathrm{B}$ memory. The asterisk ($\ast$) indicates SSC methods, for which the outputs are clustered to identify their instances, as described in Sec. 4.1.
Method MonoScene [4]* Symphonies [21]* OccFormer [63]* CGFormer [59]* IPFormer (ours)
[…]
A.6 Limitations
Experimental Results.
Experimental quantitative and qualitative results show IPFormer’s state-of-the-art performance in vision-based PSC. However, there are remaining Thing-classes (e.g. Motorcyclist) and Stuff-classes (e.g. Other-Ground) which have not been recognized, due to their low class frequency or geometric fidelity. Despite these challenges, we believe that IPFormer’s introduction of context-aware instance proposals will play a significant role in progressing 3D computer vision and specifically Panoptic Scene Completion.
Comparison with PanoSSC.
As described in Sec. 4.1, we are not able to compare the performance of IPFormer with the vision-based PSC approach of PanoSSC [47], as this method is trained on a post-processed version of SemanticKITTI that is not publicly available. We aim to collaborate with the authors of PanoSSC to train IPformer on this dataset and evaluate it on PanoSSC´s relaxed $20\text{\,}\%$ IoU threshold for matching of ground-truth and predicted instances, to present representative evaluation results. However, in Tab. 9, we provide evaluation results of IPFormer under PanoSSC’s relaxed $20\text{\,}\%$ IoU threshold. Expectedly, RQ metrics increase substantially, since more instances are recognized, while these are segmented with less fidelity. Thus, SQ metrics decrease.
Table 9: Ablation on IPFormer’s PSC and SSC performance under PanoSSC’s relaxed 2020% IoU matching threshold on the SemanticKITTI dataset.
IoU Threshold     PQ PQ-All SQ-All RQ-All PQ-Thing SQ-Thing RQ-Thing PQ-Stuff SQ-Stuff RQ-Stuff IoU     mIoU
$20\text{\,}\%$     $\mathbf{15.38}$ $\mathbf{12.74}$ $32.76$ $\mathbf{30.85}$ $\mathbf{4.31}$ $32.80$ $\mathbf{9.92}$ $\mathbf{18.88}$ $32.74$ $\mathbf{46.08}$ $\mathbf{40.90}$     $\mathbf{15.33}$
$50\text{\,}\%$     $14.45$ $6.30$ $\mathbf{41.95}$ $9.75$ $2.09$ $\mathbf{42.67}$ $3.33$ $9.35$ $\mathbf{41.43}$ $14.43$ $\mathbf{40.90}$     $\mathbf{15.33}$
A.7 Additional Ablation Experiments
Tab. 10 presents extensive additional ablation experiments on the two-stage training, the dual-head architecture, the interplay between SSC and PSC, and sensitivity to hyperparameters. All findings are in line and underscore our findings discussed in Sec. 4.
Two-Stage Training and Dual-Head Architecture.   Single-head methods (p, q) struggle to balance SSC and PSC, with single-stage configuration (p) performing the worst due to its inability to separate semantic and instance-level learning. Single-stage, dual-head methods (a, b) also fall short, as the lack of stage-wise optimization hinders instance registration. Two-stage methods highlight the advantages of stage-wise training: SSC-focused approaches (e, f, k) excel in SSC but underperform in PSC due to limited adaptation, while PSC-prioritized methods (j, n) improve instance registration but compromise semantic consistency or balance. IPFormer, method (r), decouples SSC and PSC optimization, achieving strong instance registration and balanced performance across both tasks, with a moderate SSC trade-off.
Interplay between SSC and PSC.   Across the design space, methods that emphasize SSC (e, f, k) achieve strong semantic scores but degrade PSC performance, especially PQ-Thing. Conversely, approaches prioritizing PSC (j, n, q) boost PQ-Thing or PQ-Stuff at the cost of SSC quality or overall balance. Joint or single-head variants (a, b, p, q) further struggle with instance registration or overall consistency. In contrast, our dual-head, two-stage method (r) yields the best PQ-All and strong performance across PQ-Thing and PQ-Stuff, with only a minor SSC trade-off.
Sensitivity to Hyperparameters.   We train the SSC task in Stage 1 (Eq. 8) using established hyperparameters for the cross-entropy, Semantic-SCAL, and Geometric-SCAL cost functions (Sec. 3.6). All weights associated with these functions are set to $1$, consistent with state-of-the-art and established SSC works, such as CGFormer [59], OccFormer [63], and MonoScene [4]. Nevertheless, we investigate the effect of removing the Depth loss, as its impact has not been extensively studied. Furthermore, we analyze the effect of varying hyperparameters of the Sigmoid Focal Loss [33], specifically designed to down-weight easier examples and focus the training process on harder-to-classify examples:
$\text{FL}(p)=-\alpha_{t}(1-p_{t})^{\gamma}\log(p_{t}),$ (12)
where $p_{t}$ is the predicted probability for the true class after applying the sigmoid function, $\alpha_{t}$ is the class-balancing weight, and $\gamma$ controls the down-weighting of well-classified examples.
While methods (e) and (f) in Tab. 10 achieve the highest SSC performance in terms of mIoU and IoU, respectively, their PSC performance deteriorates significantly. A similar trend is observed for method (k), which attains the second-best results for both mIoU and IoU. In contrast, the best PSC performance, measured by PQ-Thing, is achieved by method (j). Although this method demonstrates satisfactory SSC performance, it suffers from reduced PQ-Stuff and, consequently, lower PQ-All performance. Moreover, the highest PQ and PQ-Stuff performance is achieved by methods (n) and (q), respectively. However, both methods exhibit a notable decline in PQ-Thing. Specifically, method (n) experiences a substantial reduction in SSC performance, whereas method (q) maintains adequate SSC scores.
Finally, our proposed IPFormer configuration, method (r), achieves a balanced PSC performance by obtaining the best score for PQ-All and the second-best results for both PQ-Thing and PQ-Stuff, with a moderate trade-off in SSC performance.
Table 10: Further ablation experiments analyzing the sensitivity of SSC and PSC performance with respect to the hyperparameters of the objective functions, the dual-head architecture, and the two-stage training strategy. For hyperparameters, blue values indicate a difference from our proposed IPFormer configuration (r). For SSC and PSC metrics, bold and underlined values represent best and second-best results, respectively.
    Head(s)     Stage 1     Stage 2     SSC Metrics     PSC Metrics
        SSC Losses     SSC Losses     PSC Losses
        Depth CE Sem Geo     CE Sem Geo     Depth CE DICE Focal     IoU mIoU     $\text{PQ}^{{\dagger}}$ PQ-All PQ-Thing PQ-Stuff
        $\lambda_{\text{depth}}$ $\lambda_{\text{ce}}$ $\lambda_{\text{scal}}^{\text{sem}}$ $\lambda_{\text{scal}}^{\text{geo}}$     $\lambda_{\text{ce}}$ $\lambda_{\text{scal}}^{\text{sem}}$ $\lambda_{\text{scal}}^{\text{geo}}$     $\lambda_{\text{depth}}$ $\lambda_{\text{ce}}$ $\lambda_{\text{dice}}$ $\lambda_{\text{focal}}$ $\alpha$ $\gamma$
(a)     Dual         $1.00$ $1.00$ $1.00$     $0.0001$ $1.00$ $1.00$ $40.00$ $0.25$ $2.00$     $41.30$ $15.26$     $14.21$ $5.57$ $1.06$ $8.84$
(b)     Dual         $1.00$ $1.00$ $1.00$     $0.0001$ $1.00$ $1.00$ $50.00$ $0.21$ $2.30$     $39.45$ $13.06$     $14.45$ $5.03$ $0.99$ $7.79$
(c)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     $1.00$ $1.00$ $1.00$     $0.0001$ $1.00$ $1.00$ $40.00$ $0.25$ $2.00$     $43.44$ $15.79$     $14.35$ $6.06$ $1.66$ $9.27$
(d)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     $0.50$ $0.50$ $0.50$     $0.0001$ $1.00$ $1.00$ $40.00$ $0.25$ $2.00$     $43.62$ $15.30$     $14.42$ $6.27$ $1.88$ $\underline{9.46}$
(e)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     $0.20$ $0.20$ $0.20$     $0.0001$ $1.00$ $1.00$ $40.00$ $0.25$ $2.00$     $43.99$ $\mathbf{16.29}$     $14.46$ $6.05$ $1.60$ $9.29$
(f)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     $0.10$ $0.10$ $0.10$     $0.0001$ $1.00$ $1.00$ $40.00$ $0.25$ $2.00$     $\mathbf{44.13}$ $16.11$     $14.41$ $6.08$ $1.72$ $9.26$
(g)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     $0.01$ $0.01$ $0.01$     $0.0001$ $1.00$ $1.00$ $40.00$ $0.25$ $2.00$     $43.94$ $15.04$     $14.67$ $5.16$ $0.90$ $8.27$
(h)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     $1.00$ $1.00$ $1.00$     $0.0001$ $1.00$ $0.00$ $40.00$ $0.25$ $2.00$     $43.34$ $15.83$     $13.89$ $5.33$ $0.49$ $8.86$
(i)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     $0.10$ $0.10$ $0.10$     $0.0001$ $1.00$ $0.10$ $40.00$ $0.25$ $2.00$     $42.97$ $15.64$     $14.11$ $5.59$ $1.27$ $8.74$
(j)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     $0.03$ $0.03$ $0.03$     $0.0001$ $1.00$ $1.00$ $50.00$ $0.21$ $2.30$     $43.53$ $15.65$     $14.39$ $6.19$ $\mathbf{2.38}$ $8.96$
(k)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     $0.50$ $0.50$ $0.50$     $0.0001$ $1.00$ $1.00$ $50.00$ $0.21$ $2.30$     $44.07$ $16.25$     $14.23$ $5.97$ $1.67$ $9.11$
(l)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     —     $0.0001$ $1.00$ $2.00$ $40.00$ $0.25$ $2.00$     $32.45$ $7.28$     $13.92$ $5.92$ $1.58$ $9.08$
(m)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     —     $0.0001$ $1.00$ $1.00$ $50.00$ $0.21$ $2.30$     $40.45$ $14.74$     $14.59$ $5.74$ $1.53$ $8.80$
(n)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     —     $0.0001$ $1.00$ $1.00$ $45.00$ $0.21$ $2.30$     $36.73$ $11.69$     $\mathbf{15.14}$ $4.43$ $0.78$ $7.08$
(o)     Dual     $1.00$ $1.00$ $1.00$     —     $1.00$ $1.00$ $40.00$ $0.25$ $2.00$     $41.16$ $14.90$     $14.42$ $6.15$ $1.68$ $9.40$
(p)     Single     $0.0001$ $1.00$ $1.00$ $1.00$     —     $0.0001$ $1.00$ $1.00$ $40.00$ $0.25$ $2.00$     $4.90$ $2.53$     $10.66$ $0.42$ $0.13$ $0.63$
(q)     Single         —     $0.0001$ $1.00$ $1.00$ $40.00$ $0.25$ $2.00$     $42.31$ $14.80$     $14.64$ $6.23$ $1.75$ 9.49
(r)     Dual     $0.0001$ $1.00$ $1.00$ $1.00$     —     $0.0001$ $1.00$ $1.00$ $40.00$ $0.25$ $2.00$     $40.90$ $15.33$     $14.45$ 6.30 2.09 $9.35$

### v3

@@
 Existing methods, e.g.
-Symphonies symphonies , randomly initialize instance queries and incorporate context-awareness during training.
+Symphonies Jiang et al.
+(2024), randomly initialize instance queries and incorporate context-awareness during training.
 However, these queries retain their static nature at test time, as they are shared across all inputs.
 Our method IPFormer initializes them as instance proposals, which preserve context-adaptivity at test time, effectively aggregating directed features for improved PSC performance.
-Due to the the novelty of vision-based PSC and the absence of established baselines, we apply DBSCAN ester1996_dbscan clustering to Symphonies’ SSC output to retrieve its individual instances.
+Due to the the novelty of vision-based PSC and the absence of established baselines, we apply DBSCAN Ester et al.
+(1996) clustering to Symphonies’ SSC output to retrieve its individual instances.
 Panoptic Scene Completion provides a holistic scene understanding that can serve applications like autonomous driving and robotics by reconstructing volumetric geometry from sparse sensor data and assigning meaning to objects in the scene.
@@
 Despite the inherent interconnection between geometry and semantics, their separation hindered unified scene understanding.
-Classical structure-from-motion evolved into simultaneous localization and mapping (SLAM) systems for accurate reconstructions mur2015orb ; engel2014lsd ; koestler2022tandem , while deep learning facilitated monocular depth estimation (MDE) methods godard2017unsupervised ; godard2019digging ; wimbauer2021monorec , which are restricted to visible surfaces.
-In contrast, Scene Completion refers to the process of inferring the complete 3D structure of a scene, including both visible and occluded regions, from partial observations wimbauer2023behind ; han2024boosting .
-Apart from geometry, semantic image segmentation divides an image into semantically meaningful regions and labels them yu2015multi ; chen2017deeplab .
-Panoptic segmentation kirillov2019panoptic ; kirillov2019panoptic2 ; porzi2019seamless extends this approach by segmenting instances hariharan2014simultaneous for things (countable objects like person or car) while maintaining semantic segmentation for stuff (amorphous regions like road or vegetation).
+Classical structure-from-motion evolved into simultaneous localization and mapping (SLAM) systems for accurate reconstructions Mur-Artal and Tardos (2015); Engel et al.
+(2014); Koestler et al.
+(2022), while deep learning facilitated monocular depth estimation (MDE) methods Godard et al.
+(2017, 2019); Wimbauer et al.
+(2021), which are restricted to visible surfaces.
+In contrast, Scene Completion refers to the process of inferring the complete 3D structure of a scene, including both visible and occluded regions, from partial observations Wimbauer et al.
+(2023); Han et al.
+(2024).
+Apart from geometry, semantic image segmentation divides an image into semantically meaningful regions and labels them Yu and Koltun (2015); Chen et al.
+(2017).
+Panoptic segmentation Kirillov et al.
+(2019b, a); Porzi et al.
+(2019b) extends this approach by segmenting instances Hariharan et al.
+(2014) for things (countable objects like person or car) while maintaining semantic segmentation for stuff (amorphous regions like road or vegetation).
 This unified approach is vital for comprehensive scene understanding, as it distinguishes individual object instances while preserving semantic context for background elements, enabling applications like autonomous navigation.
-The recent trend in scene understanding has been to unify these tasks into Semantic Scene Completion (SSC) hayler2024s4c and Panoptic Scene Completion (PSC) pasco .
+The recent trend in scene understanding has been to unify these tasks into Semantic Scene Completion (SSC) Hayler et al.
+(2024) and Panoptic Scene Completion (PSC) Cao et al.
+(2024).
 SSC integrates geometric completion with semantic labeling, predicting both the 3D structure and semantic categories, represented as a 3D grid composed of voxels.
@@
 Semantic Scene Completion.
-The task of Semantic Scene Completion has seen a large amount of interest after it was first introduced in lmscnet .
-The initial work focused on solving the task for indoor scenes from camera data zhang2018efficient ; zhang2019cascaded ; liu2018see ; li2019depth ; li2019rgbd ; li2020anisotropic ; chen20203d ; cai2021semantic and used point clouds in outdoor scenarios with LiDAR-based methods roldao20223d ; cheng2021s3cnet ; yan2021sparse ; li2021semisupervised ; rist2021semantic .
+The task of Semantic Scene Completion has seen a large amount of interest after it was first introduced in Roldão et al.
+(2020).
+The initial work focused on solving the task for indoor scenes from camera data Zhang et al.
+(2018, 2019); Liu et al.
+(2018); Li et al.
+(2019b, a, 2020); Chen et al.
+(2020); Cai et al.
+(2021) and used point clouds in outdoor scenarios with LiDAR-based methods Roldao et al.
+(2022); Cheng et al.
+(2021); Yan et al.
+(2021); Li et al.
+(2021); Rist et al.
+(2021).
 In the following, we will discuss the most relevant contributions for Semantic Scene Completion in the context of autonomous driving.
-For a more thorough overview, we refer the reader to the surveys of xu2025survey and zhang2024vision .
-Recent progress in perception for autonomous driving has increasingly adopted Vision Transformer architectures dosovitskiyimage , notably DETR carion2020end .
+For a more thorough overview, we refer the reader to the surveys of Xu et al.
+(2025) and Zhang et al.
+(2024).
+Recent progress in perception for autonomous driving has increasingly adopted Vision Transformer architectures Dosovitskiy et al.
+(2021), notably DETR Carion et al.
+(2020).
 DETR introduced learnable object queries, which are trainable embeddings that allow the model to directly predict object sets in parallel, eliminating hand-crafted components like anchor boxes.
 In SSC, learnable queries can be adapted to predict semantic labels for 3D scene elements (e.g., voxels), streamlining the process and enhancing efficiency.
-Architectures like OccFormer occformer and VoxFormer li2023voxformer exemplify this trend, with OccFormer using multiscale voxel representations with query features to decode 3D semantic occupancy, and VoxFormer employing depth-based queries fused with image features, refined via deformable attention.
-Similarly, Symphonies symphonies refines instance queries iteratively using cross-attention between image features and 3D representations, while TPVFormer huang2023tpvformer leverages a tri-perspective view (TPV) to query scene features efficiently.
-HASSC wang2024not and CGFormer cgformer further refine this general approach by focusing on challenging voxels and contextual queries, respectively.
-Alternative approaches, such as implicit scenes via self-supervised learning in S4C, hayler2024s4c and CoHFF song2024cohff , which leverages both voxel and TPV representations to fuse features from multiple vehicles, have also emerged, though they do not utilize learnable queries.
+Architectures like OccFormer Zhang et al.
+(2023) and VoxFormer Li et al.
+(2023) exemplify this trend, with OccFormer using multiscale voxel representations with query features to decode 3D semantic occupancy, and VoxFormer employing depth-based queries fused with image features, refined via deformable attention.
+Similarly, Symphonies Jiang et al.
+(2024) refines instance queries iteratively using cross-attention between image features and 3D representations, while TPVFormer Huang et al.
+(2023) leverages a tri-perspective view (TPV) to query scene features efficiently.
+HASSC Wang et al.
+(2024a) and CGFormer Yu et al.
+(2024) further refine this general approach by focusing on challenging voxels and contextual queries, respectively.
+Alternative approaches, such as implicit scenes via self-supervised learning in S4C, Hayler et al.
+(2024) and CoHFF Song et al.
+(2024), which leverages both voxel and TPV representations to fuse features from multiple vehicles, have also emerged, though they do not utilize learnable queries.
 It is worth noting that existing literature presents ambiguity regarding context-awareness of queries for SSC.
-Although both Symphonies symphonies and CGFormer cgformer incorporate contextual cues during training, their approaches diverge at inference.
+Although both Symphonies Jiang et al.
+(2024) and CGFormer Yu et al.
+(2024) incorporate contextual cues during training, their approaches diverge at inference.
 In particular, Symphonies’ instance queries remain static and are shared across all input images, whereas CGFormer’s voxel queries preserve context-awareness by dynamically adapting to each input.
@@
 Panoptic Scene Completion.
-While panoptic LiDAR segmentation, which predicts panoptics for point clouds directly, has been widely studied Che2025S2MC-PHD ; Xiao2024lidarsegopenvoc ; xia2023lidarpanseg , voxel-based PSC approaches have only recently been introduced by PaSCo pasco .
+While panoptic LiDAR segmentation, which predicts panoptics for point clouds directly, has been widely studied Chen et al.
+(2025); Xiao et al.
+(2025); Xiang et al.
+(2023), voxel-based PSC approaches have only recently been introduced by PaSCo Cao et al.
+(2024).
 This method employs a hybrid convolutional neural network (CNN) and Transformer architecture with static instance queries to address LiDAR-based PSC, enhanced by an uncertainty-aware framework.
-In parallel, camera-based PSC from multi-frame and multi-view (surround-view) images is addressed by PanoOcc wang2024panoocc , which utilizes static voxel queries to aggregate spatiotemporal information, while SparseOcc liu2023sparseocc introduces a sparse voxel decoder and sparse static instance queries to predict semantic and instance occupancy from up to 96 temporal frames.
-Concurrently, PanoSSC shi2024panossc introduces a camera-based PSC method using forward-facing, non-temporal imagery.
+In parallel, camera-based PSC from multi-frame and multi-view (surround-view) images is addressed by PanoOcc Wang et al.
+(2024b), which utilizes static voxel queries to aggregate spatiotemporal information, while SparseOcc Liu et al.
+(2023) introduces a sparse voxel decoder and sparse static instance queries to predict semantic and instance occupancy from up to 96 temporal frames.
+Concurrently, PanoSSC Shi et al.
+(2024) introduces a camera-based PSC method using forward-facing, non-temporal imagery.
 Its dual-head TPV-based architecture separates semantic occupancy and instance segmentation tasks to infer a 3D panoptic voxel grid.
@@
 However, it overcomes a key limitation of prior approaches by avoiding reliance on static queries and instead introduces context-adaptive instance proposals that dynamically adapt to the scene at both train and test time.
-Notably, PanoSSC shi2024panossc is trained on a non-public post-processed version of the SemanticKITTI dataset semantickitti , which, in addition to the absence of released source code, challenges direct comparisons.
+Notably, PanoSSC Shi et al.
+(2024) is trained on a non-public post-processed version of the SemanticKITTI dataset Behley et al.
+(2019), which, in addition to the absence of released source code, challenges direct comparisons.
 3 Methodology
@@
 Given an input image $\mathbf{X}\in\mathbb{R}^{U\times V\times 3}$ (Fig.
-2), the backbone predicts image features $\mathbf{F}\in\mathbb{R}^{H\times W\times F}$, and the depth net predicts a depth map $\mathbf{D}\in\mathbb{R}^{U\times V\times 1}$, following VoxFormer li2023voxformer .
+2), the backbone predicts image features $\mathbf{F}\in\mathbb{R}^{H\times W\times F}$, and the depth net predicts a depth map $\mathbf{D}\in\mathbb{R}^{U\times V\times 1}$, following VoxFormer Li et al.
+(2023).
 Further refining the image features and the depth map allows for generating context-aware 3D representations, denoted as 3D context features $\mathbf{F}_{\text{3D}}$ and initial voxels $\mathbf{V}$.
@@
 In this setup, every instance proposal can eventually correlate with a single instance within the observed scene.
-The voxel proposals $\mathbf{V}_{P}$ are encoded by 3D local and global encoding, in line with CGFormer cgformer , resulting in voxel features $\mathbf{V}_{F}$.
+The voxel proposals $\mathbf{V}_{P}$ are encoded by 3D local and global encoding, in line with CGFormer Yu et al.
+(2024), resulting in voxel features $\mathbf{V}_{F}$.
 We follow a two-stage training strategy in which the first stage uses these voxel features $\mathbf{V}_{F}$ to address SSC exclusively.
@@
 3.2 3D Context Generation
-Following CGFormer cgformer , we refine the depth map $\mathbf{D}$ and the feature map $\mathbf{F}\in\mathbb{R}^{H\times W\times F}$ into per-pixel, frustum-shaped depth probabilities $\mathbf{D}_{R}\in\mathbb{R}^{H\times W\times D}$, effectively resulting in $|D|$ depth bins per pixel.
-This is achieved by applying $\Phi_{D}$ as a mixture of convolution and neighborhood cross-attention layers Hassani2023neigbor , formulated as $\mathbf{D}_{R}=\Phi_{D}(\mathbf{D},\mathbf{F}_{\text{2D}})$.
+Following CGFormer Yu et al.
+(2024), we refine the depth map $\mathbf{D}$ and the feature map $\mathbf{F}\in\mathbb{R}^{H\times W\times F}$ into per-pixel, frustum-shaped depth probabilities $\mathbf{D}_{R}\in\mathbb{R}^{H\times W\times D}$, effectively resulting in $|D|$ depth bins per pixel.
+This is achieved by applying $\Phi_{D}$ as a mixture of convolution and neighborhood cross-attention layers Hassani et al.
+(2023), formulated as $\mathbf{D}_{R}=\Phi_{D}(\mathbf{D},\mathbf{F}_{\text{2D}})$.
 Moreover, the generic image features $\mathbf{F}$ are projected into contextual features $\mathbf{F}_{\text{2D}}\in\mathbb{R}^{H\times W\times C}$ through the context net $\Phi_{C}$, denoted as $\mathbf{F}_{\text{2D}}=\Phi_{C}(\mathbf{F})$.
-The next step is to lift context features $\mathbf{F}_{\text{2D}}$ to 3D, following the design of philion2020lift .
+The next step is to lift context features $\mathbf{F}_{\text{2D}}$ to 3D, following the design of Philion and Fidler (2020).
 To elaborate, we lift them to (i) 3D context features $\mathbf{F}_{\text{3D}}$ and (ii) 3D initial voxels $\mathbf{V}$.
@@
 The summation aggregates all features that map to the same voxel, effectively accumulating evidence for stronger feature representations.
-Building upon the concept of cgformer , we add learnable embeddings to $\mathbf{V}$ that enhance its representational capacity.
+Building upon the concept of Yu et al.
+(2024), we add learnable embeddings to $\mathbf{V}$ that enhance its representational capacity.
 3.3 Proposal Initialization
@@
 Voxel Visibility.
-The initial voxels $\mathbf{V}$ are categorized into visible voxels $\mathbf{V}_{\text{vis}}$ and invisible voxels $\mathbf{V}_{\text{invis}}$, following established SSC works such as li2023voxformer .
+The initial voxels $\mathbf{V}$ are categorized into visible voxels $\mathbf{V}_{\text{vis}}$ and invisible voxels $\mathbf{V}_{\text{invis}}$, following established SSC works such as Li et al.
+(2023).
 This is achieved by first applying inverse projection $\Pi^{-1}$ on the depth map $\mathbf{D}$ with camera intrinsics $\mathbf{K}\in\mathbb{R}^{4\times 4}$.
@@
 Note that this sampling does not include learnable parameters.
-To further adapt $\mathbf{V}_{\text{vis}}$ with image context, we apply a tailored series of 3D deformable attention mechanisms, inspired by symphonies ; cgformer .
+To further adapt $\mathbf{V}_{\text{vis}}$ with image context, we apply a tailored series of 3D deformable attention mechanisms, inspired by Jiang et al.
+(2024); Yu et al.
+(2024).
 Specifically, we apply 3D deformable cross-attention (DCA) to $\mathbf{V}_{\text{vis}}$ (as queries) and the 3D context features $\mathbf{F}_{\text{3D}}$ (as keys and values), resulting in updated visible voxel features $\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}}$.
@@
 $\text{DSA}{(\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}},\tilde{\mathbf{V}}_{\text{vis}},\mathbf{x})}=\sum_{{n}={1}}^{N}A_{n}W\psi{(\tilde{\mathbf{V}}_{\text{vis}},\mathbf{x}+\Delta{\mathbf{x}})}\;.$ (4)
-Inspired by CGFormer´s cgformer query generator, our instance generator transforms the DSA output to initialize instance proposals via
+Inspired by CGFormer´s Yu et al. (2024) query generator, our instance generator transforms the DSA output to initialize instance proposals via
 $\mathbf{I}_{P}=\text{DSA}(\tilde{\mathbf{V}}_{\text{vis}}^{\mathbf{x}},\tilde{\mathbf{V}}_{\text{vis}},\mathbf{x})+\mathbf{W}_{\mathbf{I}}\in\mathbb{R}^{K\times C}\;,$ (5)
@@
 To initialize voxel proposals $\mathbf{V}_{P}$, we first merge the visible and invisible voxels through element-wise addition along dimension $C$, denoted as $\tilde{\mathbf{V}}_{\text{vis}}\oplus\mathbf{V}_{\text{invis}}$.
-We then apply 3D deformable self-attention to distribute updated context information over the entire scene volume, formulated as $\mathbf{V}_{P}=\text{DSA}(\tilde{\mathbf{V}}_{\text{vis}}\oplus\mathbf{V}_{\text{invis}})$, which is in line with li2023voxformer ; symphonies ; cgformer :
+We then apply 3D deformable self-attention to distribute updated context information over the entire scene volume, formulated as $\mathbf{V}_{P}=\text{DSA}(\tilde{\mathbf{V}}_{\text{vis}}\oplus\mathbf{V}_{\text{invis}})$, which is in line with Li et al.
+(2023); Jiang et al.
+(2024); Yu et al.
+(2024):
 3.4 Encoding
 The purpose of this part of our architecture is to (i) transform voxel proposals $\mathbf{V}_{P}$ into voxel features $\mathbf{V}_{F}$ that encode semantic information and (ii) project this semantic information onto the instance proposals $\mathbf{I}_{P}$, thus prompting semantic instance-voxel relationships.
-To encode semantic information, we propagate the voxel proposals $\mathbf{V}_{P}$ through a 3D Local and Global Encoder (LGE) based on CGFormer cgformer .
-This encoder enhances the semantic and geometric discriminability of the scene volume by integrating TPV huang2023tpvformer and voxel representations, effectively capturing global and local contextual features, respectively.
-For an in-depth discussion, we direct readers to cgformer .
+To encode semantic information, we propagate the voxel proposals $\mathbf{V}_{P}$ through a 3D Local and Global Encoder (LGE) based on CGFormer Yu et al.
+(2024).
+This encoder enhances the semantic and geometric discriminability of the scene volume by integrating TPV Huang et al.
+(2023) and voxel representations, effectively capturing global and local contextual features, respectively.
+For an in-depth discussion, we direct readers to Yu et al.
+(2024).
 The 3D LGE $\Phi_{\text{LGE}}$ transforms voxel proposals $\mathbf{V}_{P}$ into semantic voxel features $\mathbf{V}_{F}\in\mathbb{R}^{X\times Y\times Z\times C}$, formulated as $\mathbf{V}_{F}=\Phi_{\text{LGE}}(\mathbf{V}_{P})$.
@@
 In the second training stage, a lightweight panoptic head decodes instance features $\mathbf{I}_{F}$ to semantic class logits $\mathbf{I}_{L}\in\mathbb{R}^{K\times L}$, where $K$ is the number of instances and $L$ is the number of semantic classes.
-These logits are further processed with voxel features $\mathbf{V}_{F}$ to perform instance-voxel alignment that eventually enables panoptic scene aggregation, inspired by PaSCo pasco .
+These logits are further processed with voxel features $\mathbf{V}_{F}$ to perform instance-voxel alignment that eventually enables panoptic scene aggregation, inspired by PaSCo Cao et al.
+(2024).
 To elaborate, we align the instance features $\mathbf{I}_{F}$ with the voxel features $\mathbf{V}_{F}\in\mathbb{R}^{X\times Y\times Z\times C}$ to compute affinity scores $\mathbf{H}\in\mathbb{R}^{X\times Y\times Z\times K}$ by applying the dot product, resulting in $\mathbf{H}=\mathbf{V}_{F}\cdot\mathbf{I}_{F}^{\top}$.
 These scores are converted to predicted probabilities via $\mathbf{H}_{P}=\text{sigmoid}(\mathbf{H})$, where $\mathbf{H}_{P}$ denotes the model’s confidence in each voxel belonging to each instance.
-For panoptic aggregation pasco , voxel-wise logits $\mathbf{V}_{L}\in\mathbb{R}^{X\times Y\times Z\times L}$ are computed as $\mathbf{V}_{L}=\mathbf{H}_{P}\cdot\mathbf{I}_{L}$.
+For panoptic aggregation Cao et al.
+(2024), voxel-wise logits $\mathbf{V}_{L}\in\mathbb{R}^{X\times Y\times Z\times L}$ are computed as $\mathbf{V}_{L}=\mathbf{H}_{P}\cdot\mathbf{I}_{L}$.
 It should be emphasized that voxel-wise logits do not represent the probability of belonging to a certain semantic class, but the confidence of belonging to one of the instances.
@@
 $\mathcal{L}_{\text{SSC}}=\lambda_{\text{ce}}\mathcal{L}_{\text{ce}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{sem}}\mathcal{L}_{\text{scal}}^{\text{sem}}+\lambda_{\text{geo}}\mathcal{L}_{\text{scal}}^{\text{geo}}\;,$ (8)
-where the weights for the cross-entropy loss, semantic scene-class affinity loss (SCAL), and geometric SCAL monoscene are set to $\lambda_{\text{ce}}=\lambda_{\text{sem}}=\lambda_{\text{geo}}=1$, respectively, and the weight for the depth loss cgformer is set to $\lambda_{\text{depth}}=10^{-4}$.
+where the weights for the cross-entropy loss, semantic scene-class affinity loss (SCAL), and geometric SCAL Cao and de Charette (2022) are set to $\lambda_{\text{ce}}=\lambda_{\text{sem}}=\lambda_{\text{geo}}=1$, respectively, and the weight for the depth loss Yu et al.
+(2024) is set to $\lambda_{\text{depth}}=10^{-4}$.
 The second stage calculates losses mask-wise by comparing each ground-truth mask with the best matching predicted mask.
-Consistent with previous research pasco ; shi2024panossc , we find the best match by applying the Hungarian method kuhn1955hungarian to perform bipartite matching carion2020end , using an IoU threshold of $50\text{\%}$.
-This threshold follows the standard convention established in prior work on panoptic segmentation kirillov2019panoptic ; pasco , and the Hungarian algorithm ensures a globally optimal, permutation-invariant one-to-one assignment.
+Consistent with previous research Cao et al.
+(2024); Shi et al.
+(2024), we find the best match by applying the Hungarian method Kuhn (1955) to perform bipartite matching Carion et al.
+(2020), using an IoU threshold of $50\text{\%}$.
+This threshold follows the standard convention established in prior work on panoptic segmentation Kirillov et al.
+(2019b); Cao et al.
+(2024), and the Hungarian algorithm ensures a globally optimal, permutation-invariant one-to-one assignment.
 After matching, the second stage optimizes
 $\mathcal{L}_{\text{PSC}}=\lambda_{\text{mask}}\mathcal{L}_{\text{ce}}^{\text{mask}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}}+\lambda_{\text{focal}}\mathcal{L}_{\text{focal}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}\;,$ (9)
-where $\mathcal{L}_{\text{ce}}^{\text{mask}}$ represents mask-wise cross-entropy loss, $\mathcal{L}_{dice}$ is the dice loss of Milletari_2016_diceloss_vnet and $\mathcal{L}_{focal}$ is the focal loss based on Lin_2017_focalloss .
-In practice, we set the weights of these losses to $\lambda_{\text{mask}}=\lambda_{\text{dice}}=1$ and $\lambda_{\text{focal}}=40$, respectively, following PaSco pasco .
+where $\mathcal{L}_{\text{ce}}^{\text{mask}}$ represents mask-wise cross-entropy loss, $\mathcal{L}_{dice}$ is the dice loss of Milletari et al.
+(2016) and $\mathcal{L}_{focal}$ is the focal loss based on Lin et al.
+(2017).
+In practice, we set the weights of these losses to $\lambda_{\text{mask}}=\lambda_{\text{dice}}=1$ and $\lambda_{\text{focal}}=40$, respectively, following PaSco Cao et al.
+(2024).
 Moreover, $\mathcal{L}_{\text{depth}}$ is defined identically to that in $\mathcal{L}_{\text{SSC}}$.
@@
 Dataset and Baselines.
-We conduct our experiments by (i) in-domain training and testing on the SemanticKITTI SSC dataset semantickitti , and (ii) out-of-domain zero-shot generalization on the distinct SSCBench-KITTI360 li2024sscbench .
-The instance ground-truths for both datasets are provided by PaSCo pasco .
+We conduct our experiments by (i) in-domain training and testing on the SemanticKITTI SSC dataset Behley et al.
+(2019), and (ii) out-of-domain zero-shot generalization on the distinct SSCBench-KITTI360 Li et al.
+(2024).
+The instance ground-truths for both datasets are provided by PaSCo Cao et al.
+(2024).
 Given the novelty of vision-based PSC, no established baselines currently exist for direct comparison.
 Furthermore, recall from Sec.
-2 that PanoSSC shi2024panossc represents an existing vision-based PSC method using forward-facing imagery.
+2 that PanoSSC Shi et al.
+(2024) represents an existing vision-based PSC method using forward-facing imagery.
 This would make it a suitable baseline for comparison.
-However, this approach is trained on a non-public post-processed version of the SemanticKITTI dataset semantickitti , which, in addition to the absence of released source code, challenges direct comparison with our method (see Sec.
+However, this approach is trained on a non-public post-processed version of the SemanticKITTI dataset Behley et al.
+(2019), which, in addition to the absence of released source code, challenges direct comparison with our method (see Sec.
 A.6 of the technical appendix for further details).
-Consequently, we adapt the latest state-of-the-art vision-based SSC methods monoscene ; symphonies ; occformer ; cgformer and generate instance predictions by clustering Thing-classes from their outputs, in line with PaSCo pasco , which introduced the task of 3D PSC.
-In particular, to ensure fairness and methodological rigor, we apply DBSCAN ester1996_dbscan , the same Euclidean clustering algorithm used to construct the PSC ground truth in PaSCo.
+Consequently, we adapt the latest state-of-the-art vision-based SSC methods Cao and de Charette (2022); Jiang et al.
+(2024); Zhang et al.
+(2023); Yu et al.
+(2024) and generate instance predictions by clustering Thing-classes from their outputs, in line with PaSCo Cao et al.
+(2024), which introduced the task of 3D PSC.
+In particular, to ensure fairness and methodological rigor, we apply DBSCAN Ester et al.
+(1996), the same Euclidean clustering algorithm used to construct the PSC ground truth in PaSCo.
 Furthermore, we utilize pre-trained checkpoints for the baselines, acquired from the official publicly available implementations.
@@
 Out-of-Domain Zero-Shot Generalization Performance.
-Trained on SemanticKITTI semantickitti , we cross-validate our method on the distinct SSCBench-KITTI360 dataset li2024sscbench .
+Trained on SemanticKITTI Behley et al.
+(2019), we cross-validate our method on the distinct SSCBench-KITTI360 dataset Li et al.
+(2024).
 Results in Tab.
@@
 Table 1: In-domain performance on SemanticKITTI val.
-set semantickitti .
+set Behley et al.
+(2019).
 Best and second-best results are bold and underlined, respectively.
 Due to the absence of established baselines for vision-based PSC (see Sec.
-4.1), we infer state-of-the-art SSC methods and apply DBSCAN ester1996_dbscan to retrieve instances.
+4.1), we infer state-of-the-art SSC methods and apply DBSCAN Ester et al.
+(1996) to retrieve instances.
     PSC Metrics     SSC Metrics
@@
 […]
-MonoScene monoscene + DBSCAN     $10.12$ $3.43$ $15.15$ $5.33$ $0.51$ $7.36$ $0.87$ $5.56$ $20.81$ $8.57$     $36.80$ $11.31$     $\underline{4.51}$
-Symphonies symphonies + DBSCAN     $11.69$ $3.75$ $26.09$ $5.95$ $1.07$ $27.65$ $1.76$ $5.70$ $24.95$ $8.99$     $\underline{41.92}$ $15.02$     $4.54$
-OccFormer occformer + DBSCAN     $11.25$ $4.32$ $24.19$ $6.69$ $0.68$ $21.47$ $1.15$ $6.96$ $26.16$ $10.73$     $36.43$ $13.51$     $4.70$
-CGFormer cgformer + DBSCAN     $\underline{14.39}$ $\underline{6.16}$ $\mathbf{48.14}$ $\underline{9.48}$ $\mathbf{2.20}$ $\mathbf{44.46}$ $\mathbf{3.47}$ $\underline{9.03}$ $\mathbf{50.82}$ $\underline{13.86}$     $\mathbf{45.98}$ $\mathbf{16.89}$     $4.70$
+MonoScene Cao and de Charette (2022) + DBSCAN     $10.12$ $3.43$ $15.15$ $5.33$ $0.51$ $7.36$ $0.87$ $5.56$ $20.81$ $8.57$     $36.80$ $11.31$     $\underline{4.51}$
+Symphonies Jiang et al. (2024) + DBSCAN     $11.69$ $3.75$ $26.09$ $5.95$ $1.07$ $27.65$ $1.76$ $5.70$ $24.95$ $8.99$     $\underline{41.92}$ $15.02$     $4.54$
+OccFormer Zhang et al. (2023) + DBSCAN     $11.25$ $4.32$ $24.19$ $6.69$ $0.68$ $21.47$ $1.15$ $6.96$ $26.16$ $10.73$     $36.43$ $13.51$     $4.70$
+CGFormer Yu et al.
+(2024) + DBSCAN     $\underline{14.39}$ $\underline{6.16}$ $\mathbf{48.14}$ $\underline{9.48}$ $\mathbf{2.20}$ $\mathbf{44.46}$ $\mathbf{3.47}$ $\underline{9.03}$ $\mathbf{50.82}$ $\underline{13.86}$     $\mathbf{45.98}$ $\mathbf{16.89}$     $4.70$
 IPFormer (ours)     $\mathbf{14.45}$ $\mathbf{6.30}$ $\underline{41.95}$ $\mathbf{9.75}$ $\underline{2.09}$ $\underline{42.67}$ $\underline{3.33}$ $\mathbf{9.35}$ $\underline{41.43}$ $\mathbf{14.43}$     $40.90$ $\underline{15.33}$     $\mathbf{0.33}$
-Table 2: Out-of-domain zero-shot generalization performance of IPFormer and the closest baseline CGFormer+DBSCAN, by training on SemanticKITTI semantickitti and cross-validating on SSCBench-KITTI360 test set li2024sscbench .
+Table 2: Out-of-domain zero-shot generalization performance of IPFormer and the closest baseline CGFormer+DBSCAN, by training on SemanticKITTI Behley et al.
+(2019) and cross-validating on SSCBench-KITTI360 test set Li et al.
+(2024).
 IPFormer demonstrates superior absolute and relative generalization performance across PSC and SSC metrics.
@@
 SemanticKITTI
-CGFormer cgformer + DBSCAN     $14.39$ $6.16$ $\mathbf{48.14}$ $9.48$ $\mathbf{2.20}$ $\mathbf{44.46}$ $\mathbf{3.47}$ $9.03$ $\mathbf{50.82}$ $13.86$     $\mathbf{45.98}$ $\mathbf{16.89}$
+CGFormer Yu et al. (2024) + DBSCAN     $14.39$ $6.16$ $\mathbf{48.14}$ $9.48$ $\mathbf{2.20}$ $\mathbf{44.46}$ $\mathbf{3.47}$ $9.03$ $\mathbf{50.82}$ $13.86$     $\mathbf{45.98}$ $\mathbf{16.89}$
 IPFormer (ours)     $\mathbf{14.45}$ $\mathbf{6.30}$ $41.95$ $\mathbf{9.75}$ $2.09$ $42.67$ $3.33$ $\mathbf{9.35}$ $41.43$ $\mathbf{14.43}$     $40.90$ $15.33$
 KITTI-360
-CGFormer cgformer + DBSCAN     $8.44$ $1.08$ $17.82$ $1.87$ $\mathbf{0.53}$ $20.06$ $\mathbf{0.96}$ $1.48$ $16.19$ $2.54$     $28.11$ $9.44$
+CGFormer Yu et al. (2024) + DBSCAN     $8.44$ $1.08$ $17.82$ $1.87$ $\mathbf{0.53}$ $20.06$ $\mathbf{0.96}$ $1.48$ $16.19$ $2.54$     $28.11$ $9.44$
 IPFormer (ours)     $\mathbf{9.41}$ $\mathbf{1.23}$ $\mathbf{24.68}$ $\mathbf{2.16}$ $0.52$ $\mathbf{22.76}$ $0.95$ $\mathbf{1.68}$ $\mathbf{25.89}$ $\mathbf{2.93}$     $\mathbf{28.74}$ $\mathbf{9.53}$
 Relative Gap $\downarrow$
-CGFormer cgformer + DBSCAN     $41.37\%$ $82.47\%$ $62.98\%$ $80.28\%$ $75.91\%$ $54.89\%$ $72.34\%$ $83.61\%$ $68.15\%$ $81.67\%$     $38.88\%$ $44.09\%$
+CGFormer Yu et al. (2024) + DBSCAN     $41.37\%$ $82.47\%$ $62.98\%$ $80.28\%$ $75.91\%$ $54.89\%$ $72.34\%$ $83.61\%$ $68.15\%$ $81.67\%$     $38.88\%$ $44.09\%$
 IPFormer (ours)     34.88% 80.48% 41.19% 77.85% 75.12% 46.64% 71.53% 82.03% 37.52% 79.69%     29.73% 37.81%
@@
 Deep Supervision.
-Inspired by PanoSSC shi2024panossc , we investigate deep supervision in the form of auxiliary losses, essentially supervising the attention maps of each layer during the encoding of the instance proposals (Tab.
+Inspired by PanoSSC Shi et al.
+(2024), we investigate deep supervision in the form of auxiliary losses, essentially supervising the attention maps of each layer during the encoding of the instance proposals (Tab.
 5).
@@
 Figure 4: Qualitative results on the SemanticKITTI val.
-set semantickitti .
+set Behley et al.
+(2019).
 Each top row illustrates purely semantic information, following the SSC color map.